OpenStudy (anonymous):

**WILL GIVE MEDAL** Can anyone help me with this: Solve for x: -3|x + 7| = -12

OpenStudy (anonymous):

There are two answers.

OpenStudy (anonymous):

Two 'x' values.

OpenStudy (campbell_st):

start by dividing both sides of the equation by -3 |x + 7| = 4 the absolute value means there are 2 answer options... so solve x + 7 = 4 and x + 7 = -4 to get both answers...

OpenStudy (campbell_st):

then check to see if each answer for x will be a solution to the problem.